OpenStudy (tifflilly):

The line of symmetry for the quadratic equation y=ax2-8x-3isx=2 what the value of "a"

OpenStudy (vadasmommy):

a=19/4 + y/4 not 100 percent sure tho

OpenStudy (niksbhalla14mar):

y=ax^2 - 8x -3 If line of symmetry is x = 2, then in this line y is always 0. Which means for x=2 y=0 So substituting y=0 and x=2 in y=ax^2 - 8x -3 we get 0=4a-16-3 4a=19 a=19/4

OpenStudy (niksbhalla14mar):

Medal please... if you found the answer useful...

OpenStudy (loser66):

@niksbhalla14mar it is not correct. How to get the medal?

OpenStudy (niksbhalla14mar):

Solution is correct loser

OpenStudy (loser66):

I prove it is wrong!! if the answer is 19/4, then the quadratic equation is \(y=\dfrac{19}{4}x^2-8x-3\) the x-coordinate of th vertex of this parabola is \(x=-\dfrac{b}{2a}=\dfrac{16}{19}\neq 2\) Note that the line of symmetry for the quadratic must past its x -coordinate of its vertex. Hence the answer a=19/4 is NOT correct